OpenStudy (anonymous):

how do you work the problem 32x over 48y

OpenStudy (anonymous):

(2/3) * x--------->2x/3 after simplification

OpenStudy (anonymous):

actually it's 2x/3y, sorry

OpenStudy (xapproachesinfinity):

\(\huge\sf\color{maroon}{ \frac{32x}{48y} =\frac{16.2x}{16.3y} \\ simplify \ \frac{16}{16} =1}\) you get 2x/3y